OpenStudy (anonymous):

deterime if the given order pair (5/3, 1) lies on both of the lines in the system of equations by substitiuting the point of both equations. 4x+5y+5=0 and 4x=-3y

OpenStudy (anonymous):

4(5/3)+5(1)+5=0 False 4(5/3)=-3(1) False also